# Hotspot-Driven Peptide Design via Multi-Fragment Autoregressive Extension

### v1

@@ Abstract
Peptides, short chains of amino acids, interact with target proteins, making them a unique class of protein-based therapeutics for treating human diseases. Recently, deep generative models have shown great promise in peptide generation. However, several challenges remain in designing effective peptide binders. First, not all residues contribute equally to peptide-target interactions. Second, the generated peptides must adopt valid geometries due to the constraints of peptide bonds. Third, realistic tasks for peptide drug development are still lacking. To address these challenges, we introduce PepHAR, a hot-spot-driven autoregressive generative model for designing peptides targeting specific proteins. Building on the observation that certain hot spot residues have higher interaction potentials, we first use an energy-based density model to fit and sample these key residues. Next, to ensure proper peptide geometry, we autoregressively extend peptide fragments by estimating dihedral angles between residue frames. Finally, we apply an optimization process to iteratively refine fragment assembly, ensuring correct peptide structures. By combining hot spot sampling with fragment-based extension, our approach enables de novo peptide design tailored to a target protein and allows the incorporation of key hot spot residues into peptide scaffolds. Extensive experiments, including peptide design and peptide scaffold generation, demonstrate the strong potential of PepHAR in computational peptide binder design.
@@ 1 Introduction
Peptides, typically composed of $3$ to $20$ amino acid residues, are short single-chain proteins that interact with target proteins. (Bodanszky, 1988). Peptides play essential roles in various biological processes, including cellular signaling and immune responses (Petsalaki & Russell, 2008), and are emerging as a promising class of therapeutic drugs for complex diseases such as diabetes, obesity, hepatitis, and cancer (Kaspar & Reichert, 2013). Currently, there are approximately $80$ peptide drugs on the global market, $150$ in clinical development, and $400-600$ undergoing preclinical evaluation (Craik et al., 2013; Fosgerau & Hoffmann, 2015; Muttenthaler et al., 2021; Wang et al., 2022). Traditional peptide discovery methods rely on labor-intensive techniques like phage/yeast display for screening mutagenesis libraries (Boder & Wittrup, 1997; Wu et al., 2016), or energy-based computational tools to score candidate peptides (Raveh et al., 2011; Lee et al., 2018; Cao et al., 2022), both of which face limitations due to the immense combinatorial design space.
@@
Although these methods have achieved initial success in generating peptide binders with native-like structures and high affinities, several challenges remain. First, not all peptide residues contribute equally to binding. As shown in Figure 1, some residues establish key functional interactions with the target, possessing high stability and affinity. These are referred to as hot spot residues, and are critical in drug discovery (Bogan & Thorn, 1998; Keskin et al., 2005; Moreira et al., 2007). Other residues, known as scaffolds, help position the hot spots in the binding region and stabilize the peptide (Matson & Stupp, 2012; Hosseinzadeh et al., 2021). Considering the different roles of these residues, generating all of them in one step may be inefficient. Second, the generated peptides must respect the constraints imposed by peptide bonds, which are non-rotatable and enforce fixed bond lengths and planar structures (Fisher, 2001). As illustrated in Figure 1, adjacent residues must maintain specific relative positions to form proper peptide bonds. A model that represents peptide backbone structures independently as local frames (Jumper et al., 2021) may neglect these geometric constraints. Third, in practical drug discovery, peptides are not always designed from scratch. Often, initial peptide candidates are optimized, or key hot spot residues are linked via scaffold residues (Zhang et al., 2009; Yu et al., 2023). Thus, more realistic in-silico benchmarks are needed to simulate these scenarios.
To tackle these challenges, we propose PepHAR, a hot-spot-driven autoregressive generative model. By distinguishing between hot spot and scaffold residues, we break down the generation process into three stages. First, we use an energy-based density model to capture the residue distribution around the target, and apply use Langevin dynamics to sample statistically favorable and feasible hot spots. Next, instead of generating all residues simultaneously, we autoregressively extend fragments step by step, modeling dihedral angles parameterized by a von Mises distribution to maintain peptide bond geometry. Finally, since the generated fragments may not align perfectly, we apply a hybrid optimization function to assemble the fragments into a complete peptide. To simulate practical peptide drug discovery scenarios, we evaluate our method not only in de novo peptide design but also in scaffold generation, where the model scaffolds known hot spot residues into a functional peptide, akin to peptide design based on prior knowledge.
@@
  • We propose a new experimental setting, scaffold generation, to mimic practical scenarios, and demonstrate the competitive performance of our method in both peptide binder design and scaffold generation tasks.
@@ 4 Methods
To tackle the challenges in peptide design, we propose a three-stage approach to generate peptides $D$ based on their target protein $T$. Our method involves generating hot-spot residues, extending fragments, and correcting peptide structures. As shown in Figure 2 and Algorithm 1, the first stage, the Founding Stage, independently generates a small number $k$ of hot-spot residues $R_{i_{1}},...,R_{i_{k}}$ from the learned residue distribution $P(R\mid T)$. In the second stage, the Extension Stage, these hot-spot residues are used as starting points to progressively extend $k$ fragments $F_{1},...,F_{k}$ by adding new residues to the Left or Right in an autoregressive manner, until the total peptide reaches the desired length $L$. Finally, since each fragment is extended independently, the third stage, the Correction Stage, adjusts the sequences and structures of the fragments, refining them based on the gradients of the objective function to ensure valid geometries and meaningful peptide sequences.
Figure 2: Overview of our three-stage approach: In the first foudning stage, $k$ hot-spot residues are generated ($k=2$ in this example) from learned residue distribution around the target. In the second stage, new residues are extended to the fragments’ left or right based on dihedral angle distributions. Finally, in the correction stage, gradients from the objective functions are applied to refine the complete peptide.
@@ 4.1 Founding stage
The founding stage first generates $k$ hot-spot residues based on the target protein $T$. By introducing the residue distribution $P(R\mid T)$, hot-spot residues represent those that have higher probabilities (i.e., lower energies) of appearing near the binding pocket compared to other backbone structures and residue types. The generation of hot-spot residues focuses on finding regions with high probabilities, where residues are more likely to interact directly with the target. Conversely, regions with low probabilities (i.e., high energies) will have fewer peptide residues. For example, peptide residues should neither occur too far from nor too close to the pocket (Cao et al., 2022; Li et al., 2024a), as such regions may have near-zero probabilities (high energies) of residue occurrence. We parameterize $P(R\mid T)$ using an energy-based model, which is a conditional joint distribution of backbone position ${\bm{x}}$, orientation ${\bm{O}}$, and residue type $c$:
@@
here $g_{\theta}$ is a scoring function parameterized by an equivariant network, which quantifies the score of residue type $c$ occuring at a given backbone structure with position ${\bm{x}}$ and orientation ${\bm{O}}$ In other words, $g_{\theta,c}$ is the unnormalized probability of type $c$. And $Z$ is the normalizing constant associated with sequence and structure information, which we do not explicitly estimate.
@@
Training We use the Noise Contrastive Estimation (NCE) to train this parameterized energy-based model (Gutmann & Hyvärinen, 2010). NCE distinguish between samples from the true data distribution (positive points) and samples from a noise distribution (negative points). The positive distribution corresponds to the ground truth residue distribution of the peptide over the target $(c,{\bm{x}},{\bm{O}})\sim p(R\mid T)$, while the negative samples are drawn from the disturbed distribution $(c_{\mathrm{neg}},{\bm{x}}^{-},{\bm{O}}^{-})\sim p(\tilde{R}\mid T)$ by adding large spatial noises to the ground truth positions and orientations, labeled as type $c_{\mathrm{neg}}$. As positive and negative data are sampled equally, the NCE objective for a single positive data point is:
@@
As a common practice (Gutmann & Hyvärinen, 2012), we fix the negative probability $p(c_{\mathrm{neg}},{\bm{x}},{\bm{O}}\mid T)$ as a constant, simplifying the evaluation of log likelihoods for negative samples. The final loss function is given by:
@@
Since orientation lies in the $\mathrm{SO}(3)$ space, we employ the exponential map and a Riemannian random walk on the tangent space for updates (De Bortoli et al., 2022). The summation over all possible residue types ensures that we transition from regions of low occurrence probability to regions of higher probability. Finally, after each iteration, the residue type $c$ is sampled conditioned on the updated position and orientation.
@@ 4.2 Extension stage
The extension stage expands fragments into longer sequences, starting from the sampled hot-spot residues. At each extension step, we adds a new residue to either the left or right of fragment $F$. Based on the relationship between adjacent residues (Eq.1,2), the backbone structure of the new residue is inferred from its dihedral angles and the structure of the adjacent residue, which is either the N-terminal or C-terminal residue of the fragment. Specifically, when connecting a new residue to residue $R_{i_{j}}$ in the $j$ th fragment $F_{j}$, we model the dihedral angle distribution $P(\psi,\phi\mid d,R_{i_{j}},E)$, where $d\in\{\textbf{L},\textbf{R}\}$ indicates the extension direction, and $E$ represents the surrounding residues, including target $T$ and other residues in the currently generated fragments.
@@
Here, $I_{0}(\cdot)$ denotes the modified Bessel function of the first kind of order 0. The four distribution parameters are predicted by a neural network $h_{\theta}$, referred to as the prediction network. Similarly, for $d=\text{R}$, the network predicts another set of four parameters:
@@
The prediction network $h_{\theta}$ uses the same IPA backbone to extract features. However, to avoid data leakage during training, since the neighboring backbone structures are known and dihedral angles can be derived analytically, we employ directional masks in the Attention module. For instance, if the direction is Left, residues can only attend to their neighbors on the right during attention updates, and vice versa for Right.
@@ 4.3 Correction stage
For a generated peptide $D=\left[(c_{1},{\bm{x}}_{1},{\bm{O}}_{1}),...,(c_{N},{\bm{x}}_{N},{\bm{O}}_{N})\right]$, the dihedrals of each residue are derived based on the backbone structures of adjacent residues. To ensure self-consistency between dihedrals and backbone structures, we use the dihedrals to estimate new backbone structures and compare them with the original ones. The distance between these backbone structures reflects the validity of the generated peptide with respect to peptide bond properties and planarity. We define the distance between two residues’ backbone structures separately for position and orientation, and derive the backbone objective considering both directions:
@@
The final optimization objective is a weighted sum of the backbone and dihedral objectives. We iteratively update the peptide’s backbone structures by taking gradients, similar to the founding stage, but here we optimize the entire peptide at each timestep. The residue types are predicted by the density model $g_{\theta}$ at the end of each update step. Unlike the founding stage, where we started from random structures, the correction stage begins with the complete peptide.
@@ 5.1 Peptide Binder Design
Baselines We compare PepHAR against three state-of-the-art peptide design models. RFDiffusion (Watson et al., 2022) uses pre-trained weights from RoseTTAFold (Baek et al., 2021) and generates protein backbone structures through a denoising diffusion process. Peptide sequences are then recovered using ProteinMPNN (Dauparas et al., 2022). ProteinGenerator augument RFDiffusion with sequence-structure jointly generation (Lisanza et al., 2023). PepFlow (Li et al., 2024a) models full-atom peptide and samples peptides using a flow-matching framework on a Riemannian manifold. PepGLAD (Kong et al., 2024) employs equivariant latent diffusion networks to generate full-atom peptide structures.
Results As shown in Table  1, PepHAR effectively generates peptides that exhibit valid geometries, native-like structures, and improved energies. While RFDiffusion produces valid peptides due to its pretrained protein folding weights, PepFlow, which is trained solely on peptide datasets, struggles with generating valid peptides, making it challenging for practical applications. In contrast, PepHAR’s autoregressive generation based on dihedral angles not only ensures the production of valid peptides but also allows for precise placement at the binding site with accurate secondary structures. Similar to previous work (Li et al., 2024a), RFDiffusion excels at generating stable peptide-target structures, while PepHAR demonstrates competitive performance compared to PepFlow. Additionally, PepHAR shows impressive results in terms of novelty and diversity, highlighting its potential for exploring peptide distributions and designing a wide range of peptides for real-world applications. Figure 4 illustrates two examples of peptides generated by PepHAR, which closely resemble the structures and binding sites of native peptides while exhibiting low binding energies, indicating high affinities for the target.
@@ 5.2 Peptide Scaffold Generation
Baselines and Metrics We use the same baselines and metrics as in the Peptide Design task. Specifically, for RFDiffusion and Protein Generator, the known hot spot residues are provided as an additional condition, along with the target. For PepFlow, we modify the ODE sampling process by initializing it with the ground truth hot spot residues and ensure the model to only modify the remaining residues. In our method, we replace the sampled hot spot residues with the known ground truth residue.
Results As shown in Table  2, PepHAR demonstrates excellent performance in scaffolding hot spot residues into complete peptides. Given that the hot spot residues have functional binding capabilities, while the scaffold residues contribute primarily to structural integrity, the generated peptides are expected to possess valid, native structures with high stability. PepHAR successfully generates valid and native-like structures, ensuring that scaffold residues do not disrupt interactions between hot spot residues, achieving the best scores in SSR and BSR. Moreover, PepHAR achieves competitive stability results compared to RFDiffusion, which is trained on a larger PDB dataset. Additionally, PepHAR produces novel and diverse scaffolds. Figure 5 presents two examples of scaffolded peptides alongside native peptides and given residues. The generated scaffolds exhibit similar structures to the native ones, while displaying variations in geometry and orientation at the midpoints and ends of the peptides, indicating flexibility in the scaffolding regions. Furthermore, the generated scaffolds often have lower total energy than the native peptides, suggesting enhanced stability of the complex and improved interaction potential.

### v2

@@ Abstract
Peptides, short chains of amino acids, interact with target proteins, making them a unique class of protein-based therapeutics for treating human diseases. Recently, deep generative models have shown great promise in peptide generation. However, several challenges remain in designing effective peptide binders. First, not all residues contribute equally to peptide-target interactions. Second, the generated peptides must adopt valid geometries due to the constraints of peptide bonds. Third, realistic tasks for peptide drug development are still lacking. To address these challenges, we introduce PepHAR, a hot-spot-driven autoregressive generative model for designing peptides targeting specific proteins. Building on the observation that certain hot spot residues have higher interaction potentials, we first use an energy-based density model to fit and sample these key residues. Next, to ensure proper peptide geometry, we autoregressively extend peptide fragments by estimating dihedral angles between residue frames. Finally, we apply an optimization process to iteratively refine fragment assembly, ensuring correct peptide structures. By combining hot spot sampling with fragment-based extension, our approach enables de novo peptide design tailored to a target protein and allows the incorporation of key hot spot residues into peptide scaffolds. Extensive experiments, including peptide design and peptide scaffold generation, demonstrate the strong potential of PepHAR in computational peptide binder design. Source code will be available at https://github.com/Ced3-han/PepHAR.
@@ 1 Introduction
Peptides, typically composed of $3$ to $20$ amino acid residues, are short single-chain proteins interacting with target proteins. (Bodanszky, 1988). Peptides play essential roles in various biological processes, including cellular signaling and immune responses (Petsalaki & Russell, 2008). They are emerging as a promising class of therapeutic drugs for complex diseases such as diabetes, obesity, hepatitis, and cancer (Kaspar & Reichert, 2013). Currently, there are approximately $80$ peptide drugs on the global market, $150$ in clinical development, and $400-600$ undergoing preclinical evaluation (Craik et al., 2013; Fosgerau & Hoffmann, 2015; Muttenthaler et al., 2021; Wang et al., 2022). Traditional peptide discovery methods rely on labor-intensive techniques like phage/yeast display for screening mutagenesis libraries (Boder & Wittrup, 1997; Wu et al., 2016), or energy-based computational tools to score candidate peptides (Raveh et al., 2011; Lee et al., 2018; Cao et al., 2022), both of which face limitations due to the immense combinatorial design space.
@@
Although these methods have initially succeeded in generating peptide binders with native-like structures and high affinities, several challenges remain. First, not all peptide residues contribute equally to binding. As shown in Figure 1, some residues establish key functional interactions with the target, possessing high stability and affinity. These are referred to as hot spot residues and are critical in drug discovery (Bogan & Thorn, 1998; Keskin et al., 2005; Moreira et al., 2007). Other residues, known as scaffolds, help position the hot spots in the binding region and stabilize the peptide (Matson & Stupp, 2012; Hosseinzadeh et al., 2021). Considering the different roles of these residues, generating all of them in one step may be inefficient. Second, the generated peptides must respect the non-rotatable constraints imposed by peptide bonds, which enforce fixed bond lengths and planar structures (Fisher, 2001). As illustrated in Figure 1, adjacent residues must maintain specific relative positions to form proper peptide bonds. A model that represents peptide backbone structures independently as local frames (Jumper et al., 2021) may neglect these geometric constraints. Third, peptides are not always designed from scratch in practical drug discovery. Initial peptide candidates are often optimized, or key hot spot residues are linked via scaffold residues (Zhang et al., 2009; Yu et al., 2023). Thus, more realistic in-silico benchmarks are needed to simulate these scenarios.
To tackle these challenges, we propose PepHAR, a hot-spot-driven autoregressive generative model. By distinguishing between hot spot and scaffold residues, we break down the generation process into three stages. First, we use an energy-based density model to capture the residue distribution around the target and apply Langevin dynamics to sample statistically favorable and feasible hot spots. Next, instead of generating all residues simultaneously, we autoregressively extend fragments step by step, modeling dihedral angles parameterized by a von Mises distribution to maintain peptide bond geometry. Finally, since the generated fragments may not align perfectly, we apply a hybrid optimization function to assemble the fragments into a complete peptide. To simulate practical peptide drug discovery scenarios, we evaluate our method not only in de novo peptide design but also in scaffold generation, where the model scaffolds known hot spot residues into a functional peptide, akin to peptide design based on prior knowledge.
@@
  • We propose a new experimental setting, scaffold generation, to mimic practical scenarios and demonstrate the competitive performance of our method in both peptide binder design and scaffold generation tasks.
@@ 4 Methods
To tackle the challenges in peptide design, we propose a three-stage approach to generate peptides $D$ based on their target protein $T$. Our method generates hot-spot residues, extends fragments, and corrects peptide structures. As shown in Figure 2 and Algorithm 1, the first stage, the Founding Stage, independently generates a small number $k$ of hot-spot residues $R_{i_{1}},...,R_{i_{k}}$ from the learned residue distribution $P(R\mid T)$. In the second stage, the Extension Stage, these hot-spot residues are used as starting points to progressively extend $k$ fragments $F_{1},...,F_{k}$ by adding new residues to the Left or Right in an autoregressive manner, until the total peptide reaches the desired length $L$. Finally, since each fragment is extended independently, the third stage, the Correction Stage, adjusts the sequences and structures of the fragments, refining them based on the gradients of the objective function to ensure valid geometries and meaningful peptide sequences.
Figure 2: Overview of our three-stage approach: In the first foudning stage, $k$ hot-spot residues are generated ($k=2$ in this example) from learned residue distribution around the target. New residues are extended to the fragments’ left or right in the second stage based on dihedral angle distributions. Finally, in the correction stage, gradients from the objective functions are applied to refine the complete peptide.
@@ 4.1 Founding stage
The founding stage first generates $k$ hot-spot residues based on the target protein $T$. By introducing the residue distribution $P(R\mid T)$, hot-spot residues represent those with higher probabilities (i.e., lower energies) of appearing near the binding pocket compared to other backbone structures and residue types. The generation of hot-spot residues focuses on finding regions with high probabilities, where residues are more likely to interact directly with the target. Conversely, regions with low probabilities (i.e., high energies) will have fewer peptide residues. For example, peptide residues should neither occur far from nor close to the pocket (Cao et al., 2022; Li et al., 2024a), as such regions may have near-zero probabilities (high energies) of residue occurrence. We parameterize $P(R\mid T)$ using an energy-based model, which is a conditional joint distribution of backbone position ${\bm{x}}$, orientation ${\bm{O}}$, and residue type $c$:
@@
here $g_{\theta}$ is a scoring function parameterized by an equivariant network, which quantifies the score of residue type $c$ occurring at a given backbone structure with position ${\bm{x}}$ and orientation ${\bm{O}}$ In other words, $g_{\theta,c}$ is the unnormalized probability of type $c$. And $Z$ is the normalizing constant associated with sequence and structure information, which we do not explicitly estimate.
@@
Training We use the Noise Contrastive Estimation (NCE) to train this parameterized energy-based model (Gutmann & Hyvärinen, 2010). NCE distinguishes between samples from the true data distribution (positive points) and samples from a noise distribution (negative points). The positive distribution corresponds to the ground truth residue distribution of the peptide over the target $(c,{\bm{x}},{\bm{O}})\sim p(R\mid T)$, while the negative samples are drawn from the disturbed distribution $(c_{\mathrm{neg}},{\bm{x}}^{-},{\bm{O}}^{-})\sim p(\tilde{R}\mid T)$ by adding large spatial noises to the ground truth positions and orientations, labeled as type $c_{\mathrm{neg}}$. As positive and negative data are sampled equally, the NCE objective for a single positive data point is:
@@
As a common practice (Gutmann & Hyvärinen, 2012), we fix the negative probability $p(c_{\mathrm{neg}},{\bm{x}},{\bm{O}}\mid T)$ as a constant, simplifying the evaluation of log-likelihoods for negative samples. The final loss function is given by:
@@
Since orientation lies in the $\mathrm{SO}(3)$ space, we employ the exponential map and a Riemannian random walk on the tangent space for updates (De Bortoli et al., 2022). The summation of all possible residue types ensures that we transition from regions of low occurrence probability to regions of higher probability. Finally, after each iteration, the residue type $c$ is sampled conditioned on the updated position and orientation.
@@ 4.2 Extension stage
The extension stage expands fragments into longer sequences, starting from the sampled hot-spot residues. At each extension step, we add a new residue to either the left or right of fragment $F$. Based on the relationship between adjacent residues (Eq.1,2), the backbone structure of the new residue is inferred from its dihedral angles and the structure of the adjacent residue, which is either the N-terminal or C-terminal residue of the fragment. Specifically, when connecting a new residue to residue $R_{i_{j}}$ in the $j$ th fragment $F_{j}$, we model the dihedral angle distribution $P(\psi,\phi\mid d,R_{i_{j}},E)$, where $d\in\{\textbf{L},\textbf{R}\}$ indicates the extension direction and $E$ represents the surrounding residues, including target $T$ and other residues in the currently generated fragments.
@@
Here, $I_{0}(\cdot)$ denotes the modified Bessel function of the first kind of order 0. The four distribution parameters are predicted by a neural network $h_{\theta}$, called the prediction network. Similarly, for $d=\text{R}$, the network predicts another set of four parameters:
@@
The prediction network $h_{\theta}$ uses the same IPA backbone to extract features. However, to avoid data leakage during training, we employ directional masks in the Attention module since the neighboring backbone structures are known and dihedral angles can be derived analytically. For instance, if the direction is Left, residues can only attend to their neighbors on the right during attention updates, and vice versa for Right.
@@ 4.3 Correction stage
For a generated peptide $D=\left[(c_{1},{\bm{x}}_{1},{\bm{O}}_{1}),...,(c_{N},{\bm{x}}_{N},{\bm{O}}_{N})\right]$, the dihedrals of each residue are derived based on the backbone structures of adjacent residues. To ensure self-consistency between dihedrals and backbone structures, we use them to estimate new backbone structures and compare them with the original ones. The distance between these backbone structures reflects the validity of the generated peptide with respect to peptide bond properties and planarity. We define the distance between two residues’ backbone structures separately for position and orientation and derive the backbone objective considering both directions:
@@
The final optimization objective is a weighted sum of the backbone and dihedral objectives. We iteratively update the peptide’s backbone structures by taking gradients, similar to the founding stage, but we optimize the entire peptide at each timestep. The density model $g_{\theta}$ predicts the residue types at the end of each update step. Unlike the founding stage, where we started from random structures, the correction stage begins with the complete peptide.
@@ 5.1 Peptide Binder Design
Baselines We compare PepHAR against three state-of-the-art peptide design models. RFDiffusion (Watson et al., 2022) uses pre-trained weights from RoseTTAFold (Baek et al., 2021) and generates protein backbone structures through a denoising diffusion process. Peptide sequences are then recovered using ProteinMPNN (Dauparas et al., 2022). ProteinGenerator augment RFDiffusion with sequence-structure jointly generation (Lisanza et al., 2023). PepFlow (Li et al., 2024a) models full-atom peptides and samples peptides using a flow-matching framework on a Riemannian manifold. PepGLAD (Kong et al., 2024) employs equivariant latent diffusion networks to generate full-atom peptide structures.
Results As shown in Table  1, PepHAR effectively generates peptides that exhibit valid geometries, native-like structures, and improved energies. While RFDiffusion produces valid peptides due to its pre-trained protein folding weights, PepFlow, which is trained solely on peptide datasets, struggles with generating valid peptides, making it challenging for practical applications. In contrast, PepHAR’s autoregressive generation based on dihedral angles ensures the production of valid peptides and allows for precise placement at the binding site with accurate secondary structures. Similar to previous work (Li et al., 2024a), RFDiffusion excels at generating stable peptide-target structures, while PepHAR demonstrates competitive performance compared to PepFlow. Additionally, PepHAR shows impressive results in terms of novelty and diversity, highlighting its potential for exploring peptide distributions and designing a wide range of peptides for real-world applications. Figure 4 illustrates two examples of peptides generated by PepHAR, which closely resemble the structures and binding sites of native peptides while exhibiting low binding energies, indicating high affinities for the target.
@@ 5.2 Peptide Scaffold Generation
Baselines and Metrics We use the same baselines and metrics as in the Peptide Design task. Specifically, for RFDiffusion and Protein Generator, the known hot spot residues, along with the target, are provided as an additional condition. For PepFlow, we modify the ODE sampling process by initializing it with the ground truth hot spot residues and ensuring the model only modifies the remaining residues. In our method, we replace the sampled hot spot residues with the known ground truth residue.
Results As shown in Table  2, PepHAR demonstrates excellent performance in scaffolding hot spot residues into complete peptides. Given that the hot spot residues have functional binding capabilities. In contrast, the scaffold residues contribute primarily to structural integrity, the generated peptides are expected to possess valid, native structures with high stability. PepHAR successfully generates valid and native-like structures, ensuring that scaffold residues do not disrupt interactions between hot spot residues, achieving the best scores in SSR and BSR. Moreover, PepHAR achieves competitive stability compared to RFDiffusion, which is trained on a larger PDB dataset. Additionally, PepHAR produces novel and diverse scaffolds. Figure 5 presents two examples of scaffolded peptides alongside native peptides and given residues. The generated scaffolds exhibit similar structures to the native ones, while displaying variations in geometry and orientation at the midpoints and ends of the peptides, indicating flexibility in the scaffolding regions. Furthermore, the generated scaffolds often have lower total energy than the native peptides, suggesting enhanced stability of the complex and improved interaction potential.